Pattern of OLED Issues

There are a number of threads discussing various issues with the OLED…

Fan Speed Disables/Corrupts OLED
OLED Fading
OLED Display Malfunction
OLED Fades to Black
OLED Not Working After Update
LCD (OLED) Display Just Went Almost Black

Since receiving my EON, I’ve had several of the problems mentioned in the threads above. The latest experience I had yesterday occurred following a fresh install of Bullseye. I installed the script and then configured the fan, OLED, and RTC. I created a RAID1 via mdadm and left it initializing overnight. This morning, the array was finished initializing but the OLED display was completely blank. I tried the following steps to bring the display back (note that the OLED cable is secure on both ends and was not touched during any of the steps below)…

  • Cycled argononed.service (no change).
  • Restart of the Pi (no change).
  • Shutdown via the power switch and then power on (no change).
  • Shutdown via the power switch, pull power cord, and then power on (top two lines of OLED displayed).
  • Shutdown via the power switch, pull power cord, let sit for a few hours, and then power on (OLED back to normal).

The threads above suggest a disproportionate number of similar problems for ~1000 units. It appears (to me) that there is some underlying issue with the OLED. Whether it’s the OLED itself, the script driving it, or another hardware issue with the EON, it’s difficult to say.

I’m hoping Argon 40 will weigh in and make some recommendations of how to handle/troubleshoot these issues further to help identify the root cause and a solution.


1 Like

Based on a cursory review of the scripting that handles the OLED, I’m suspecting there is likely problems with how the buffer is being cleared and written to for the items belong displayed. It may require a different way of clearing the buffer or hitting it with the same command twice.

I haven’t had time to dig into it as I’d like to-but the garbled display seems like a problem with clearing the buffer properly and then things go out of whack. It is the clearing of the buffer by completely powering off completely that is what “fixes” the garble issue.

Now, as for it not working in an OS, that is likely due to a missing dependency. You might be able to diagnose that based on logs.

1 Like

Just wanted to throw in my experience. I had the above happen to me yesterday after a reboot, left it unplugged overnight, then it was fine when booting up today.
Otherwise, the OLED had functioned perfectly for weeks up until this point.

My display went dark again after five days of uptime.

I suspect the code is not properly communicating with the OLED which results in (over time) a failure to communicate. This in turn requires a power down for long enough to drain what ever caps might be in the path… Just a feeling though.